The radius of a piece of 22-gauge tungsten wire is 0.321 mm and the resistivity of tungsten is 1.5 x 10-8 Ω m. Calculate the resistance per unit length of the wire and if a potential difference of 1.5V is maintained across a tungsten wire of length 2.00 m, what is the current in the wire?


1
Expert's answer
2021-03-21T11:38:36-0400

"r=0.321\ mm=0.321\cdot10^{-3}\ m"


"\rho=1.5\cdot10^{-8}\ \Omega\cdot m"


"V=1.5\ V"


"l=1\ m"


"L=2\ m"


"R, I-?"



"R_0=\rho\frac{l}{A}=\rho\cdot \frac{l}{\pi r^2}=1.5\cdot10^{-8}\cdot \frac{1}{3.14\cdot (0.321\cdot10^{-3})^2}=0.046\ (\Omega)" . Answer


"I=\frac{V}{R}=\frac{V}{2R_0}=\frac{1.5}{2\cdot0.046}=16.3\ (A)" . Answer